OpenStudy (anonymous):

i have a homework assignment and it just states_prime factor a.13 b.18 c.36 d.41 e.344 f.10,200----what does he want me to do?

OpenStudy (anonymous):

it wants you take factor all the numbers into products of primes

OpenStudy (anonymous):

13x1

OpenStudy (anonymous):

a) 13 is already prime, done b) 18=2*9=2*3*3 prime factors are 2 and 3 c)36=6*6=2*2*3*3 prime factors are 2 and 3

OpenStudy (anonymous):

thanks so much